OpenStudy (anonymous):

calculate the integral of the following using:

OpenStudy (anonymous):

the standard form \[\int\limits_{}^{}f'(x)e ^{f(x)}dx=e ^{f(x)}+c , or \int\limits_{}^{}e ^{u}(du/dx)dx = e ^{u}+c\]

OpenStudy (anonymous):

\[(a) 2xe ^{x ^{2}+3}\] \[(b)(10x-2)e ^{5x ^{2}-2x}\] \[(c)(3x+2)e ^{3x ^{2}+4x+1}\] \[(d)(x ^{2}-2x)e ^{x ^{3}-3x ^{2}}\]

OpenStudy (watchmath):

What have you try with these problems?

OpenStudy (anonymous):

yea i have i can do (a),(b) but c and d i got the wrong answers

OpenStudy (watchmath):

You need to adjust a little bit for c) and d). For example in c) if you use \(u=3x^2+4x\) then \(du=6x+4=2(3x+2)\, dx\). So your integral become \(\frac{1}{2}\int e^u\, du\)

OpenStudy (anonymous):

oh i understand thanks but what happens to the (3x+2) inside the bracket

OpenStudy (watchmath):

\((3x+2)\,dx\) become \(\frac{1}{2}du\)

OpenStudy (anonymous):

ok so how about this one \[(6x ^{2}-8x+6)e ^{x ^{3}-2x ^{2}+3x-5}\]

OpenStudy (watchmath):

what is the derivative of \(x^3-2x^2+3x-5\) ?

OpenStudy (anonymous):

3x^2-4x+3

OpenStudy (watchmath):

Good! How that derivative relates to \(6x^2-8x+6\) ?

OpenStudy (anonymous):

its half of it right

OpenStudy (anonymous):

the coefficients are doubled

OpenStudy (watchmath):

Yes great! Then for \(f(x)=x^3-2x^2+3x-5\) the integral is \(\frac{1}{2}\int f'(x)e^{f(x)}\,dx\)

OpenStudy (watchmath):

Sorry your were right :)

OpenStudy (watchmath):

So \(2\int f'(x)e^{f(x)}\,dx\).

OpenStudy (anonymous):

oh i see how it works out now lol:) ok thanks ill have shot at the others thankyou

OpenStudy (watchmath):

Great! :D